# thanks: Corresponding authors:

### v1

@@ III.1.1 Regimes of bistability
We observe that even under the idealized assumption that the cell could freely choose its effector concentration, bistable behavior arises only within a restricted range of parameter values. This is shown numerically in the red regions of Fig. 11. To investigate the conditions under which multiple steady states are possible, we also derive analytical bounds in parameter space in Appendix  B.1. By setting $d\bar{A}/d\bar{t}=0$ and rewriting the resulting expression in standard polynomial form, we can infer the number of positive real roots using classical results (i.e., Descartes’ rule of signs) that relate sign changes in the coefficients of a polynomial to the number of positive roots. This leads to a necessary condition on the parameters of the system: for bistability to be possible for at least one value of the effector concentration, the system must satisfy the inequalities
@@
Figure 12: Comparison between Hill function and thermodynamic treatments of bistability in the auto-activation switch. (A) The Hill function model predicts bistability even when it is forbidden in the thermodynamic formulation by the Descartes bounds. Shared parameters are $\bar{r}_{0}=0.1$, $\bar{r}_{2}=2$. Two illustrative thermodynamic curves are shown here in blue as the two remaining parameters, $\bar{r}_{1}$ and $\omega$, are varied. Regardless of the choices of $\bar{r}_{1}$ and $\omega$, the thermodynamic curves exhibit no bistability in the bistable region of the Hill function. One thermodynamic curve shown here have parameters $\bar{r}_{1}=0.1$, $\omega=2$; the other has parameters $\bar{r}_{1}=1$, $\omega=100$. (B) Hill function fails to capture the tunability of the EC50 in the thermodynamics formulation. By changing the parameters $\bar{r}_{1}$ and $\omega$, the switch from high to low concentrations of $\bar{A}$ can be tuned to occur at different effector concentration $c$. Shared parameters are $\bar{r}_{0}=0.1$, $\bar{r}_{2}=10$. The left thermodynamic curve shown here has parameters $\bar{r}_{1}=1$, $\omega=1$. The left thermodynamic curve shown here has parameters $\bar{r}_{1}=1$, $\omega=50$. Nevertheless, all Hill function curves switch at the same effector concentration, regardless of Hill coefficient $n$, with the evidence shown in Fig. 31 in Appendix F. The Hill functions shown in (A) and (B) have $n=2$.
@@
We note that the necessary criteria for bistability derived in Eqns. 1012 do not impose any effective constraint when tuning either $\bar{r}_{0}$, $\bar{r}_{1}$, or $\bar{r}_{2}$ for the set of parameters chosen in Fig. 11. Specifically, the predicted lower bound on $\bar{r}_{1}$ exceeds the upper limit on $\bar{r}_{1}$ allowed by the auto-activation constraint (i.e., $\bar{r}_{1}\leq\bar{r}_{2}$). Similarly, the threshold for $\bar{r}_{2}$ above which bistability is possible lies below $\bar{r}_{1}$, meaning the system no longer behaves as a strictly auto-activating unit. Indeed, we observe that for parameter values consistent with the auto-activation regime, when either $\bar{r}_{1}$ or $\bar{r}_{2}$ is varied individually in Fig. 11(C) and (D), the system fails to exhibit bistability only in a narrow region where $\bar{r}_{1}$ approaches $\bar{r}_{2}$. Note, however, that this behavior is not universal. For different parameter values, the relative positioning of these thresholds may change, and the necessary conditions specified by Eqns.  10-12 could become more explicitly informative. The conclusions drawn here are therefore specific to the parameter set used in this analysis.
@@
Interestingly, we observe bistability for values of the cooperativity parameter $\omega$ that are less than one, as seen in Fig. 11(A). This can be reconciled in two complementary ways. First, we can define an effective cooperativity for the system, given by $\omega_{\text{eff}}=\omega\cdot\bar{r}_{2}/2$. From the necessary conditions for bistability derived in Eqns. 10 and  12 we find that $\omega_{\text{eff}}>1$ is required for bistability. While this condition is necessary but not sufficient, it suggests that $\omega_{\text{eff}}$ captures the functional cooperativity of the system more accurately than $\omega$ alone, as it incorporates both the interaction between bound activators and the maximal rate of activator production. We can also reconcile our bistable results containing values of $\omega$ less than one by examining the effective Hill coefficient of the steady-state input–output function. As shown in Appendix G, bistability is observed only when the effective Hill coefficient exceeds one. This reinforces the idea that the system can exhibit bistability even when $\omega<1$, provided the overall nonlinearity—quantified either by $\omega_{\text{eff}}$ or the effective Hill coefficient—is sufficiently strong.
@@ III.1.2 Comparing Hill function and thermodynamic formulations
As Hill himself tells us, this functional form provides a summary of the occupancy of hemoglobin (the example he used, though it has been applied much more broadly). If we think of the huge topic of input-output functions in biology, then the kind of characteristics embodied in the Hill approach include a representation of leakiness (the amount of output even in the absence of input, $p_{\text{bound}}(0)$), dynamic range, $EC_{\text{50}}$ (the input concentration at which the output reaches half its maximum, $EC_{\text{50}}=K$) and the sensitivity as measured by the slope of the input-output curve (usually in logarithmic variables) at the midpoint. It is instructive to hear Hill commenting on his thinking: “My object was rather to see whether an equation of this type can satisfy all the observations, than to base any direct physical meaning on $n$ and $K$ [62].” He goes further in his 1913 paper noting [63] “In point of fact $n$ does not turn out to be a whole number, but this is due simply to the fact that aggregation is not into one particular type of molecule, but rather into a whole series of different molecules: so that equation (1) (our Eqn. 13) is a rough mathematical expression for the sum of several similar quantities with $n$ equal to 1, 2, 3, 4 and possibly higher integers.” We think it important to remember that the Hill function is a phenomenological description of equilibrium binding that assumes certain states of occupancy are irrelevant, or alternatively, that bunches all of the states of occupancy into one effective non-integer state of occupancy.
@@
In comparing and contrasting the two treatments of transcription factor binding, we find that they can yield different results. Hill functions approximate away some thermodynamic details (such as cooperativity) and describe the dynamics as
$\begin{split}\frac{d\bar{A}}{d\bar{t}}=-\bar{A}&+\frac{\bar{r}_{0}+\bar{r}_{2}(p_{\text{act}}(c)\bar{A})^{n}}{1+(p_{\text{act}}(c)\bar{A})^{n}},\end{split}$ (14)
where $n$ is the Hill coefficient. We observe that the functional form of the production term is similar to that of the thermodynamic model. In Appendix F, we further show that the probabilities of each state of transcription factor binding are similar between the Hill function and the thermodynamic model. Nevertheless, the steady states predicted by Hill functions surprisingly differ from the thermodynamics prediction in two principal ways.
First, Hill functions can contradict the thermodynamic model in their prediction of bistability, as shown in Fig. 12(A). We compute bifurcation curves—reporting the fixed points of the system—using both Hill and thermodynamic formulations under identical baseline parameters $\bar{r}_{0}=0.1$ and $\bar{r}_{2}=2$. For the thermodynamic model, we additionally explore a wide range of intermediate activation strengths $\bar{r}_{1}\in[0.1,2]$ (ensuring $\bar{r}_{0}\leq\bar{r}_{1}\leq\bar{r}_{2}$) and cooperativity values $\omega\in[10^{-4},10^{4}]$. The Hill function model with Hill coefficient $n=2$ predicts a bistable regime at low $c$, highlighted in light orange. In contrast, across all combinations of $\bar{r}_{1}$ and $\omega$, the thermodynamic model exhibits no bistability in this same region, demonstrating that the Hill formulation can introduce bistability not permitted by physical thermodynamic constraints.
Second, the Hill function model does not capture the variability in the effector concentration at which the switch is flipped—that is, the critical $c$ value at which the system transitions from a high to a low steady-state level of $A$. In Fig. 12(B), we again compute the bifurcation curves with $\bar{r}_{0}=0.1$ and $\bar{r}_{2}=10$ for both the thermodynamic and the Hill model. The thermodynamic model predicts that this threshold depends sensitively on both $\bar{r}_{1}$ and $\omega$, allowing the switch-like transition to occur over a broad range of effector concentrations. In contrast, all Hill function curves, regardless of Hill coefficient, undergo the transition at approximately the same $c$ value, shown by the orange circle in Fig. 12(B). We show Hill function bifurcation curves with other Hill coefficients in Appendix F to demonstrate this fact. From this point of view, the thermodynamic model predicts a more flexible switch than the Hill function model. Given the relatively light computational requirements for the model systems discussed here, as opposed to more complex and high-dimensional gene-interaction networks [64], we will proceed with the thermodynamic formulation for all other gene circuits considered in the paper.
@@ III.1.3 Timescale for stabilization
Figure 13: Dynamics of the auto-activation switch. The parameters of the system are fixed at: $\omega=7.5$, $\bar{r}_{0}=0.1$, $\bar{r}_{1}=1$, $\bar{r}_{2}=20$, and $c=2\cdot 10^{-5}\text{ M}$. (A) Time evolution of the activator concentration for various initial conditions. Each black curve represents a trajectory $\bar{A}(\bar{t})$ starting from a different initial condition $\bar{A}_{0}$. Dashed horizontal lines indicate the stable ($\bar{A}_{\mbox{high}}$ and $\bar{A}_{\mbox{low}}$) and unstable ($\bar{A}_{\mbox{unstable}}$) fixed points. (B) Relaxation timescale obtained from exponential fits to the trajectory $\bar{A}(\bar{t})$ as a function of the initial concentration $\bar{A}_{0}$. Vertical lines indicate the positions of the steady states, while horizontal dashed lines mark the timescales associated with small perturbations around each fixed point, computed from the inverse slope $|f^{\prime}(\bar{A})|^{-1}$ of the function $f$ defined in Eqn. 16. The solid horizontal black line corresponds to the reference timescale $\bar{t}=1$.
@@
As shown in Fig. 13(A), the relaxation timescale depends strongly on the initial activator concentration $\bar{A}_{0}$. In particular, as seen in the figure, the relaxation timescale to steady state increases as the initial value of activator concentration, $\bar{A}_{0}$, gets closer to the unstable fixed point. The increased time to reach steady state near the unstable point reflects the typical behavior of positive feedback systems near bifurcations, where the system lingers near the threshold before switching  [65]. However, for large initial concentrations of $A$, the dimensionless relaxation timescale approaches 1, as shown by the line at $\bar{t}=1$ in Fig. 13(A). This corresponds to the protein degradation timescale $1/\gamma$, which serves as the unit of time in our nondimensionalization. This also corresponds to the relaxation timescale of a simply regulated gene, activated solely by an upstream transcription factor. We find that auto-activation consistently delays the response compared to simple regulation, although the two converge in the high-$A$ limit. Within the Hill-function framework, positive autoregulation has likewise been shown to slow gene circuit response times [66], in contrast to the accelerating effect of negative feedback [67].
@@
for sufficiently small $\delta\bar{A}(\bar{t})$. The function $f$ is given by
@@
These insights are reflected in Fig. 13(B), where we numerically compute the relaxation timescale as a function of the initial condition $\bar{A}_{0}$. For initial values near the stable fixed points, the timescale closely follows the linear prediction $|f^{\prime}(\bar{A}_{\text{low/high}})|^{-1}$. In contrast, when the initial condition lies near the unstable fixed point, the system initially diverges exponentially and only later relaxes nonlinearly to a stable state. This leads to an overall increase in the time to reach steady state, consistent with the trajectories displayed in Fig. 13(A) (see Appendix H for the precise numerical procedure used to compute these timescales). Together, these results reveal that the stabilization timescale in auto-activation circuits is not fixed but depends sensitively on initial conditions, particularly near unstable fixed points—highlighting the importance of considering nonlinear transient dynamics when predicting response times in bistable gene networks.
@@ III.2 The mutual repression regulatory motif
where we have also defined an effector with concentration $c_{1}$ that regulates the activity of repressor 1. The production term attributes a rate $r$ to the state with no bound repressors as shown in Fig. 14. As in auto-activation, this analysis only accounts for the presence of RNA polymerase implicitly, which we discuss further in Appendix A.2.
We can write dimensionless forms of these kinetic equations by transforming $\bar{t}=\gamma t$, $\bar{r}_{i}=R_{i}/K_{2}$, and $\bar{r}=r/\gamma K_{2}$, and obtain
@@
Figure 17: Bistability regimes in mutual repression as a function of the relative DNA-binding affinities of repressors $R_{1}$ and $R_{2}$. Colored regions denote distinct bistable phase space geometries, defined by whether bistability occurs for a finite window of $c_{1}$, a finite window of $c_{2}$, or the presence of both at sufficiently small concentrations. Otherwise, the system is never bistable for any concentration $(c_{1},c_{2})$ in the interval $[10^{-7}\text{M},10^{-4}\text{M}]$ considered. These geometries evolve as the binding affinity ratio $\bar{K}=K_{1}/K_{2}$ is varied, with $\omega_{1}=\omega_{2}=7.5$ and $\bar{r}=2$ held constant.
@@ III.2.1 Conditions for bistability
The first geometry (marked orange in the legend of Fig. 17) corresponds to a situation where bistability is present only when $c_{1}$ lies within a finite interval $[c_{1}^{\min},c_{1}^{\max}]$ for a given $c_{2}$. In this case, both bounds of the interval are strictly positive and finite, and $c_{2}$ must be smaller than a certain threshold. The limiting factor is therefore $c_{1}$, which must be finely tuned to enable bistability, while $c_{2}$ simply needs to remain below a critical level. Nevertheless, when decreasing $c_{2}$ the bistable interval in $c_{1}$ broadens, showing that lower $c_{2}$ expands the range of $c_{1}$ values supporting bistability.
@@
In both cases, we observe the same sequence of transitions in the structure of the bistable region, as shown in Appendix  I.3. High values of $\omega_{2}$ reflect strong cooperative binding of repressor $R_{2}$ to the DNA, meaning that binding becomes more favorable when two repressors are present. This effect mirrors what happens when increasing $\bar{K}$: if $\bar{K}>1$, then $K_{1}>K_{2}$, implying that higher concentrations of $R_{1}$ are required for effective DNA binding compared to $R_{2}$. As a result, increasing $\bar{K}$ effectively enhances the influence of $R_{2}$, analogous to increasing $\omega_{2}$. Conversely, tuning $\omega_{1}$ affects the system similarly, but with the roles of $c_{1}$ and $c_{2}$ reversed. The effects of cooperativity are examined in greater detail in Appendix I.3.
Figure 18: Bistability regions in mutual repression as a function of cooperativity and production rate. In both panels, the two repressors have equal DNA-binding affinities, with $\bar{K}=1$. We determine the boundary in $(\omega_{1},\omega_{2})$ space that separates regions where bistability is possible for some $(c_{1},c_{2})$ from those where it is impossible. (A) Example for $\bar{r}=1$. (B) Rather than showing the full bistability map, we report only this boundary curve for increasing $\bar{r}$.
From Fig. 17, we note that extreme (low or high) values of $\bar{K}$ tend to suppress bistability, as they strongly favor one repressor over the other across all concentrations. In contrast, high values of $\omega_{1}$ or $\omega_{2}$ amplify repression mostly when the corresponding repressor is present at high concentration. As a result, the system requires finely tuned inducer concentrations to counteract this cooperative imbalance and sustain bistability, effectively constraining the range of inducer concentrations for which the system can be bistable—as shown in Appendix I.3.
@@
In Fig. 18(A), we classify parameter combinations in the $(\omega_{1},\omega_{2})$ plane for $\bar{r}=1$ according to whether the system exhibits bistability for any inducer concentrations $(c_{1},c_{2})$. Fig. 18(B) shows how the boundary between monostable and bistable regimes shifts with $\bar{r}$, separating regions where bistability is either inaccessible or achievable for at least some inducer pairs. We observe that the production rate, when coupled to cooperativity, plays a critical role in enabling bistability—much like in the auto-activation system, where the product $\omega\bar{r}_{2}$ must exceed a threshold to generate bistability. We quantify this cooperative relationship between $\omega_{i}$ and $\bar{r}$ further in Appendix I.1 by deriving a necessary condition for bistability where
@@
In auto-activation systems, the effective Hill coefficient can vary above or below one, but bistability only occurs when it exceeds one. In contrast, for mutual repression, the effective Hill coefficients for $R_{1}$ and $R_{2}$ vary with parameters but remain strictly greater than one when $\omega_{1},\omega_{2}>0$ (Appendix I.1), making the Hill coefficient less informative about the existence of bistability. Even in models using empirical Hill functions to describe the production terms, Hill coefficients above one are not always sufficient for bistability, and computational studies show that extended network interactions can yield bistability even with coefficients below one [70].

### v2

@@ III.1.1 Regimes of bistability
We observe that even under the idealized assumption that the cell could freely choose its effector concentration, bistable behavior arises only within a restricted range of parameter values. This is shown numerically in the red regions of Fig. 11. To investigate the conditions under which multiple steady states are possible, we also derive analytical bounds in parameter space in Appendix B.1. By setting $d\bar{A}/d\bar{t}=0$ and rewriting the resulting expression in standard polynomial form, we can infer the number of positive real roots using classical results (i.e., Descartes’ rule of signs) that relate sign changes in the coefficients of a polynomial to the number of positive roots. This leads to a necessary condition on the parameters of the system: for bistability to be possible for at least one value of the effector concentration, the system must satisfy the inequalities
@@
We note that the necessary criteria for bistability derived in Eqns. 10 - 12 do not impose any effective constraint when tuning either $\bar{r}_{0}$, $\bar{r}_{1}$, or $\bar{r}_{2}$ for the set of parameters chosen in Fig. 11. Specifically, the predicted lower bound on $\bar{r}_{1}$ exceeds the upper limit on $\bar{r}_{1}$ allowed by the auto-activation constraint (i.e., $\bar{r}_{1}\leq\bar{r}_{2}$). Similarly, the threshold for $\bar{r}_{2}$ above which bistability is possible lies below $\bar{r}_{1}$, meaning the system no longer behaves as a strictly auto-activating unit. Indeed, we observe that for parameter values consistent with the auto-activation regime, when either $\bar{r}_{1}$ or $\bar{r}_{2}$ is varied individually in Fig. 11(C) and (D), the system fails to exhibit bistability only in a narrow region where $\bar{r}_{1}$ approaches $\bar{r}_{2}$. Note, however, that this behavior is not universal. For different parameter values, the relative positioning of these thresholds may change, and the necessary conditions specified by Eqns. 10 - 12 could become more explicitly informative. The conclusions drawn here are therefore specific to the parameter set used in this analysis.
@@
Interestingly, we observe bistability for values of the cooperativity parameter $\omega$ that are less than one, as seen in Fig. 11(A). This can be reconciled in two complementary ways. First, we can define an effective cooperativity for the system, given by $\omega_{\text{eff}}=\omega\cdot\bar{r}_{2}/2$. From the necessary conditions for bistability derived in Eqns. 10 and 12 we find that $\omega_{\text{eff}}>1$ is required for bistability. While this condition is necessary but not sufficient, it suggests that $\omega_{\text{eff}}$ captures the functional cooperativity of the system more accurately than $\omega$ alone, as it incorporates both the interaction between bound activators and the maximal rate of activator production. We can also reconcile our bistable results containing values of $\omega$ less than one by examining the effective Hill coefficient of the steady-state input–output function. As shown in Appendix H, bistability is observed only when the effective Hill coefficient exceeds one, as anticipated from theoretical considerations [32]. This reinforces the idea that the system can exhibit bistability even when $\omega<1$, provided the overall nonlinearity—quantified either by $\omega_{\text{eff}}$ or the effective Hill coefficient—is sufficiently strong.
@@ III.1.2 Comparing Hill function and thermodynamic formulations
As Hill himself tells us, this functional form provides a summary of the occupancy of hemoglobin (the example he used, though it has been applied much more broadly). If we think of the huge topic of input-output functions in biology, then the kind of characteristics embodied in the Hill approach include a representation of leakiness (the amount of output even in the absence of input, $p_{\text{bound}}(0)$), dynamic range, $EC_{\text{50}}$ (the input concentration at which the output reaches half its maximum, $EC_{\text{50}}=K$) and the sensitivity as measured by the slope of the input-output curve (usually in logarithmic variables) at the midpoint. It is instructive to hear Hill commenting on his thinking: “My object was rather to see whether an equation of this type can satisfy all the observations, than to base any direct physical meaning on $n$ and $K$ [37].” He goes further in his 1913 paper noting [38] “In point of fact $n$ does not turn out to be a whole number, but this is due simply to the fact that aggregation is not into one particular type of molecule, but rather into a whole series of different molecules: so that equation (1) (our Eqn. 13) is a rough mathematical expression for the sum of several similar quantities with $n$ equal to 1, 2, 3, 4 and possibly higher integers.” We think it is important to remember that the Hill function is a phenomenological description of equilibrium binding that assumes certain states of occupancy are irrelevant, or alternatively, that bunches all of the states of occupancy into one effective non-integer state of occupancy [90].
@@
In comparing and contrasting the two treatments of transcription factor binding, we find that they can yield different results. Hill functions approximate away some thermodynamic states and describe the dynamics as
$\begin{split}\frac{dA}{dt}=-\gamma A&+\frac{r_{0}+r_{2}(p_{\text{act}}(c)\frac{A}{K_{d}^{\text{eff}}})^{n}}{1+(p_{\text{act}}(c)\frac{A}{K_{d}^{\text{eff}}})^{n}},\end{split}$ (14)
where $n$ is the Hill coefficient and $K_{d}^{\text{eff}}$ is an effective dissociation constant. Appendix F demonstrates explicitly how this Hill function form derives from the high cooperativity limit of the thermodynamic model, removing the single activator-bound state as has been reported previously in the literature [90].
[…]
Reconciling the Hill and thermodynamic models for $n=2$ in the limit of large $\omega$ requires an effective dissociation constant of $K_{d}^{\text{eff}}=K_{d}/\sqrt{\omega}$, as discussed in Appendix F. Since this means that $K_{d}^{\text{eff}}\neq K_{d}$ when $\omega\neq 1$, the natural concentration scales differ in the Hill and thermodynamic settings, and we must thus exercise caution in comparison. The remaining analysis of this subsection therefore employs the dimensional forms of each model to allow direct comparison.
When cooperativity $\omega$ is small or moderate, the Hill function model can provide different predictions from the thermodynamic model. There are two ways to compare the models. One compares only the probability of state occupancy, as done in Appendix F, and the other compares the bifurcation curves of the full dynamical system when each production term is used. In the first case, the two models often yield similar end results, with their state occupancy probabilities agreeing to some degree even for small $\omega$. In contrast, when the production terms are embedded in the dynamical system, the models can diverge substantially. Figure 12(A) shows this for the low cooperativity case with $\omega=1$. Most notably, at low effector concentrations the Hill model shows bistability while the thermodynamic model does not. In the intermediate cooperativity regime of Fig. 12(B), the bifurcation diagrams resemble each other more closely. Nevertheless, they still make conflicting predictions for bistability in a range of effector concentration between $10^{-6}$ M to $10^{-5}$ M. In the high cooperativity regime of Fig. 12(C), the two models ultimately agree and produce identical bifurcation profiles.
Given the relatively light computational requirements for the model systems discussed here, as opposed to more complex and high-dimensional gene-interaction networks [79, 80], we will proceed with the thermodynamic formulation at high cooperativity for all other gene circuits considered in the paper. Since our analysis addresses the existence of bistability across all three cooperativity regimes, using the full thermodynamic model also ensures that the discussion captures all essential dynamical features across varying cooperativities.
@@ III.1.3 Timescale for stabilization
Figure 13: Dynamics of the auto-activation switch. The parameters of the system are fixed at: $\omega=7.5$, $\bar{r}_{0}=0.1$, $\bar{r}_{1}=1$, $\bar{r}_{2}=20$, and $c=2\cdot 10^{-5}\text{ M}$. (A) Time evolution of the activator concentration for various initial conditions. Each black curve represents a trajectory $\bar{A}(\bar{t})$ starting from a different initial condition $\bar{A}_{0}$. Dashed horizontal lines indicate the stable ($\bar{A}_{\rm high}$ and $\bar{A}_{\rm low}$) and unstable ($\bar{A}_{\rm unstable}$) fixed points. (B) Relaxation timescale obtained from exponential fits to the trajectory $\bar{A}(\bar{t})$ as a function of the initial concentration $\bar{A}_{0}$. Vertical lines indicate the positions of the steady states, while horizontal dashed lines mark the timescales associated with small perturbations around each fixed point, computed from the inverse slope $|f^{\prime}(\bar{A})|^{-1}$ of the function $f$ defined in Eqn. 16. The solid horizontal black line corresponds to the reference timescale $\bar{t}=1$.
@@
As shown in Fig. 13(A), the relaxation timescale depends strongly on the initial activator concentration $\bar{A}_{0}$. In particular, as seen in the figure, the relaxation timescale to steady state increases as the initial value of activator concentration, $\bar{A}_{0}$, gets closer to the unstable fixed point. The increased time to reach steady state near the unstable point reflects the typical behavior of positive feedback systems near bifurcations, where the system lingers near the threshold before switching [97]. However, for large initial concentrations of $A$, the dimensionless relaxation timescale approaches 1, as shown by the line at $\bar{t}=1$ in Fig. 13(A). This corresponds to the protein degradation timescale $1/\gamma$, which serves as the unit of time in our non-dimensionalization. This also corresponds to the relaxation timescale of a simply regulated gene, activated solely by an upstream transcription factor. We find that auto-activation consistently delays the response compared to simple regulation, although the two converge in the high-$A$ limit. Within the Hill-function framework, positive autoregulation has likewise been shown to slow gene circuit response times [60], in contrast to the accelerating effect of negative feedback [78].
@@
for sufficiently small $\delta\bar{A}(\bar{t})$, where $f^{\prime}(\bar{A})$ is a derivative with respect to $\bar{A}$ rather than a time derivative. The function $f$ is given by
@@
These insights are reflected in Fig. 13(B), where we numerically compute the relaxation timescale as a function of the initial condition $\bar{A}_{0}$. For initial values near the stable fixed points, the timescale closely follows the linear prediction $|f^{\prime}(\bar{A}_{\text{low/high}})|^{-1}$. In contrast, when the initial condition lies near the unstable fixed point, the system initially diverges exponentially and only later relaxes nonlinearly to a stable state. This leads to an overall increase in the time to reach steady state, consistent with the trajectories displayed in Fig. 13(A) (see Appendix I for the precise numerical procedure used to compute these timescales). Together, these results reveal that the stabilization timescale in auto-activation circuits is not fixed but depends sensitively on initial conditions, particularly near unstable fixed points—highlighting the importance of considering nonlinear transient dynamics when predicting response times in bistable gene networks.
@@ III.2 The mutual repression regulatory motif
where we have also defined an effector with concentration $c_{1}$ that regulates the activity of repressor 1. The production term attributes a rate $r$ to the state with no bound repressors as shown in Fig. 14. As in auto-activation, this analysis only accounts for the presence of RNA polymerase implicitly, which we discuss further in Appendix A.2. Note that having two effectors means that the control knob becomes a 2-dimensional vector $(c_{1},c_{2})$. The bistable region becomes an area, and $(c_{1},c_{2})$ can be tuned following arbitrary trajectories in the inducer concentration space. We will explore these threads in this section.
To continue the analysis, we can write dimensionless forms of these kinetic equations by transforming $\bar{t}=\gamma t$, $\bar{R}_{i}=R_{i}/K_{2}$, and $\bar{r}=r/\gamma K_{2}$, and obtain
@@
Figure 17: Bistability regimes in mutual repression as a function of the relative DNA-binding affinities of repressors $R_{1}$ and $R_{2}$. Colored regions denote distinct bistable phase space geometries, defined by whether bistability occurs for a finite window of $c_{1}$, a finite window of $c_{2}$, or the presence of both at sufficiently small concentrations. Otherwise, the system is never bistable for any concentration $(c_{1},c_{2})$ in the interval $10^{-7}\text{M},10^{-4}\text{M}]$ considered. These geometries evolve as the binding affinity ratio $\bar{K}=K_{1}/K_{2}$ is varied, with $\omega_{1}=\omega_{2}=7.5$ and $\bar{r}=2$ held constant.
@@ III.2.1 Conditions for bistability
The first geometry (marked orange in the legend of Fig. 17) corresponds to a situation where bistability is present only when $c_{1}$ lies within a finite interval $c_{1}^{\min},c_{1}^{\max}]$ for a given $c_{2}$. In this case, both bounds of the interval are strictly positive and finite, and $c_{2}$ must be smaller than a certain threshold. The limiting factor is therefore $c_{1}$, which must be finely tuned to enable bistability, while $c_{2}$ simply needs to remain below a critical level. Nevertheless, when decreasing $c_{2}$ the bistable interval in $c_{1}$ broadens, showing that lower $c_{2}$ expands the range of $c_{1}$ values supporting bistability.
@@
In both cases, we observe the same sequence of transitions in the structure of the bistable region, as shown in Appendix J.3. High values of $\omega_{2}$ reflect strong cooperative binding of repressor $R_{2}$ to the DNA, meaning that binding becomes more favorable when two repressors are present. This effect mirrors what happens when increasing $\bar{K}$: if $\bar{K}>1$, then $K_{1}>K_{2}$, implying that higher concentrations of $R_{1}$ are required for effective DNA binding compared to $R_{2}$. As a result, increasing $\bar{K}$ effectively enhances the influence of $R_{2}$, analogous to increasing $\omega_{2}$. Conversely, tuning $\omega_{1}$ affects the system similarly, but with the roles of $c_{1}$ and $c_{2}$ reversed. The effects of cooperativity are examined in greater detail in Appendix J.3.
Figure 18: Bistability regions in mutual repression as a function of cooperativity and production rate. In both panels, the two repressors have equal DNA-binding affinities, with $\bar{K}=1$. We determine the boundary in $(\omega_{1},\omega_{2})$ space that separates regions where bistability is possible for some $(c_{1},c_{2})$ from those where it is not possible. (A) Example for $\bar{r}=1$. (B) Boundary curves in the $(\omega_{1},\omega_{2})$ plane showing how the region of possible bistability expands as $\bar{r}$ increases. Each curve corresponds to fixed $\bar{r}$ values and separates parameter combinations that are monostable for all $(c_{1},c_{2})$ (below the curve) from combinations that are bistable for some $(c_{1},c_{2})$ (above the curve). Colors indicate the value of $\bar{r}$ used to compute each boundary curve.
From Fig. 17, we note that extreme (low or high) values of $\bar{K}$ tend to suppress bistability, as they strongly favor one repressor over the other across all concentrations. In contrast, high values of $\omega_{1}$ or $\omega_{2}$ amplify repression mostly when the corresponding repressor is present at high concentration. As a result, the system requires finely tuned inducer concentrations to counteract this cooperative imbalance and sustain bistability, effectively constraining the range of inducer concentrations for which the system can be bistable—as shown in Appendix J.3.
@@
In Fig. 18(A), we classify parameter combinations in the $(\omega_{1},\omega_{2})$ plane for $\bar{r}=1$ according to whether the system exhibits bistability for any inducer concentrations $(c_{1},c_{2})$. Fig. 18(B) shows how the boundary between monostable and bistable regimes shifts with $\bar{r}$, separating regions where bistability is either inaccessible or achievable for at least some inducer pairs. We observe that the production rate, when coupled to cooperativity, plays a critical role in enabling bistability—much like in the auto-activation system, where the product $\omega\bar{r}_{2}$ must exceed a threshold to generate bistability. We quantify this cooperative relationship between $\omega_{i}$ and $\bar{r}$ further in Appendix J.1 by deriving a necessary condition for bistability where
@@
In auto-activation systems, the effective Hill coefficient can vary above or below one, but bistability only occurs when it exceeds one. In contrast, for mutual repression, the effective Hill coefficients for $R_{1}$ and $R_{2}$ vary with parameters but remain strictly greater than one when $\omega_{1},\omega_{2}>0$ (Appendix J.1), making the Hill coefficient less informative about the existence of bistability. Even in models using empirical Hill functions to describe the production terms, Hill coefficients above one are not always sufficient for bistability, and computational studies show that extended network interactions can yield bistability even with coefficients below one [71].